# STI-Bench: Are MLLMs Ready for Precise Spatial-Temporal World Understanding?

### v3

@@ Acknowledgements
Appendix A Technical Details for Benchmark Construction and Analysis
@@ A.2 Ground Truth Computation
In STI-Bench, the ground truth annotations are derived from multi-modal data, including frame-by-frame camera intrinsics/extrinsics and precise 3D point cloud annotations. We utilize three publicly available datasets (Omni6DPose, ScanNet, Waymo) to cover different scenarios (desktop, indoor, outdoor). Below, we provide eight benchmark tasks with unified notations and formulas.
@@ A.3 Human Involved Quality Control
Despite leveraging automated methods for generating large quantities of question-answer pairs, several critical issues necessitated human intervention for quality assurance. These included inaccuracies or ambiguities in automatically generated object descriptions, logical inconsistencies or irrelevant content within questions and options, and insufficient visual information in video data due to occlusions or limited resolution. To address these challenges, we established a dedicated online platform for systematic multi-round manual review and random sampling checks. Human reviewers evaluated each QA pair based on several criteria: clarity and accuracy of object descriptions, logical coherence between questions and answer options, sufficiency of visual information to accurately respond to questions, and effectiveness of distractors in posing genuine challenges. After passing the rigorous review process, answer options were randomized to prevent bias arising from fixed option ordering. Ultimately, this meticulous procedure enabled us to select over 2,000 high-quality QA pairs from more than 300 video sequences, significantly enhancing the robustness and reliability of the STI-Bench benchmark dataset.
Appendix B Evaluation Details
The evaluation framework of STI-Bench is carefully designed to assess multimodal models across diverse real-world scenarios, including Desktop, Indoor, and Outdoor environments. Specifically, the benchmark incorporates tasks derived from three publicly available datasets: Omni6DPose for desktop-scale, ScanNet for indoor, and Waymo for outdoor. Each task comprises question-answer pairs generated using Multimodal Large Language Models, which provide detailed semantic descriptions for objects based on video frames and corresponding multimodal data.
To ensure task diversity and challenge, each question is accompanied by carefully constructed distractors, differing significantly yet plausibly from the ground truth. Recognizing the variance in precision requirements among different scenarios—millimeter-level for Desktop, centimeter-to-decimeter-level for Indoor, and decimeter-to-meter-level for Outdoor—we apply scenario-specific scaling factors to generate distractors. Logarithmic sampling techniques are also utilized to ensure an even distribution of numerical differences, thus maintaining evaluation robustness.
Furthermore, rigorous human quality control was conducted, involving multiple rounds of manual filtering and random sampling reviews. This step ensured the logical consistency, clarity, and accuracy of all QA pairs. Finally, answer options were randomized post-review to prevent ordering bias. As a result, the final benchmark contains more than 2,000 meticulously curated QA pairs, providing a comprehensive and reliable evaluation framework across diverse real-world scenarios.
Appendix C Visualization Results
To qualitatively illustrate the effectiveness and interpretability of STI-Bench, visualization examples from representative Desktop, Indoor, and Outdoor scenarios are presented. Visualizations include video frames with annotated bounding boxes and corresponding object descriptions generated by Multimodal Large Language Models (MLLMs). For instance, objects are annotated with clear descriptions such as “a beige minivan with a roof rack” or “a red backpack on a brown leather sofa,” which are directly visible in the annotated frames. Additionally, visual representations of object trajectories, camera movements, and bounding box projections are provided, clearly demonstrating the spatial relationships and dynamic interactions within each scenario. These visualizations not only validate the accuracy and robustness of generated annotations but also provide intuitive insights into the model’s interpretability and performance in diverse real-world settings. Additionally, we also showcase typical examples of the Gemini 2.5 Pro’s thought process as used during the error analysis.

### v4

@@ Acknowledgements
Appendix A Details for Benchmark Construction
@@ A.2 Ground Truth Computation
In STI-Bench, the ground truth annotations are derived from multi-modal data, including frame-by-frame camera intrinsics/extrinsics and precise 3D point cloud annotations. We utilize three publicly available datasets (Waymo, ScanNet, Omni6DPose) to cover different scenarios (outdoor, indoor, desktop). Below, we provide eight benchmark tasks with unified notations and formulas.
@@ A.3 Human Involved Quality Control
Despite leveraging automated methods for generating large quantities of question-answer pairs, several critical issues necessitated human intervention for quality assurance. These included inaccuracies or ambiguities in automatically generated object descriptions, logical inconsistencies or irrelevant content within questions and options, and insufficient visual information in video data due to occlusions or limited resolution. To address these challenges, we established a dedicated online platform for systematic multi-round manual review and random sampling checks. Human reviewers evaluated each QA pair based on several criteria: clarity and accuracy of object descriptions, logical coherence between questions and answer options, sufficiency of visual information to accurately respond to questions, and effectiveness of distractors in posing genuine challenges. After passing the rigorous review process, answer options were randomized to prevent bias arising from fixed option ordering. Ultimately, approximately 20% of all generated QA pairs were retained. This meticulous procedure enabled us to select over 2,000 high-quality QA pairs from more than 300 video sequences, significantly enhancing the robustness and reliability of the STI-Bench benchmark dataset.
Appendix B Adjustment Details
To ensure option diversity under different positioning accuracies, we generate refined distractors for each scene $S\in\{\mathrm{desktop},\mathrm{indoor},\mathrm{outdoor}\}$. Given the ground-truth answer $A_{\text{c}}$ and four initial distractors $\{A_{d,1},A_{d,2},A_{d,3},A_{d,4}\}$, the adjustment consists of three steps.
Step 1: Scene-dependent error interval.
Define $[E_{\min},E_{\max}]$ as
[…]
Step 2: Log-uniform sampling of target error.
Sample $u\sim\mathcal{U}(0,1)$ and compute
$e=E_{\min}\Bigl{(}\tfrac{E_{\max}}{E_{\min}}\Bigr{)}^{u}.$ (10)
Step 3: Joint adjustment of distractors.
All distractors are shifted towards $A_{\text{c}}$ with a common weight $w$:
$A_{d,i}^{\prime}=(1-w)A_{d,i}+wA_{\text{c}},\qquad i=1,\dots,4.$ (11)
The weight $w$ is chosen such that the closest adjusted distractor attains the target error $e$:
$w=1-\frac{e}{\displaystyle\min_{i}\bigl{\lVert}A_{d,i}-A_{\text{c}}\bigr{\rVert}}.$ (12)
By sharing the same $w$, the minimum gap between any distractor and the correct answer is exactly $e$, while others remain further apart, thereby preserving option dispersion.
Appendix C Detailed Results
To demonstrate the effectiveness of STI-Bench, we present visualization examples of QA pairs from representative outdoor (see Figure 8), indoor (see Figure 9), and desktop (see Figure 10) scenarios.
To provide a representative analysis of the capability bottlenecks of current SOTA models, we sampled over 200 erroneous reasoning processes from Gemini 2.5 Pro and presented one typical correct and one typical incorrect example (see Figure 11 and Figure 12).
In addition, to more concretely quantify the performance of different models on fine-grained categories, Table 4, Table 5, and Table 6 provide a detailed breakdown of model performance across various fine-grained categories within outdoor, indoor, and desktop scenarios, respectively.
@@ Appendix C Detailed Results
Table 4: Performance on Fine-grained Outdoor Categories.
  Category (Outdoor) Models Avg.
Gemini- 2.5-Pro Qwen2.5-VL -72B Claude-3.7 -Sonnet InternVL 2.5-78B
  Night 37.2% 43.6% 38.5% 35.9% 38.8%
Day 50.3% 47.9% 44.7% 43.3% 46.6%
Dawn/Dusk 45.3% 50.4% 47.0% 47.4% 47.5%
Rainy 51.1% 55.5% 54.5% 48.6% 52.4%
Clear 51.0% 59.4% 50.0% 55.7% 54.0%
Foggy 71.4% 28.6% 71.4% 71.4% 60.7%
Table 5: Performance on Fine-grained Indoor Categories.
  Category (Indoor) Models Avg.
Gemini- 2.5-Pro Qwen2.5-VL -72B Claude-3.7 -Sonnet InternVL 2.5-78B
  laundry room 25.0% 25.0% 50.0% 50.0% 37.5%
copy/mail room 27.6% 30.0% 30.0% 40.0% 31.9%
bookstore/library 32.0% 24.0% 32.0% 28.0% 29.0%
kitchen 26.7% 29.0% 32.3% 32.3% 30.1%
bedroom/hotel 36.6% 33.9% 42.3% 30.4% 35.8%
office 36.2% 33.3% 37.4% 30.2% 34.3%
hallway 38.5% 42.0% 38.5% 42.0% 40.3%
living room/lounge 39.0% 34.6% 40.3% 29.9% 36.0%
conference room 26.8% 33.9% 38.6% 32.2% 32.9%
apartment 39.3% 48.3% 46.4% 44.8% 44.7%
misc 36.4% 45.5% 54.5% 45.5% 45.5%
bathroom 49.1% 34.2% 34.2% 41.2% 39.7%
storage/basement/garage 30.8% 57.1% 35.7% 57.1% 45.2%
Table 6: Performance on Fine-grained Desktop Categories.
  Category (Desktop) Models Avg.
Gemini- 2.5-Pro Qwen2.5-VL -72B Claude-3.7 -Sonnet InternVL 2.5-78B
  Stone 27.3% 16.7% 33.3% 9.1% 21.6%
Paper 40.0% 37.5% 18.8% 25.0% 30.3%
Leather 23.1% 33.3% 40.0% 26.7% 30.8%
Plastic 32.8% 32.6% 28.6% 29.8% 31.0%
Fabric 40.7% 35.5% 30.0% 35.5% 35.4%
Wood 37.8% 33.7% 34.9% 33.7% 35.0%
Tile 39.6% 38.0% 38.0% 40.0% 38.9%
Concrete 42.9% 33.3% 53.3% 46.7% 44.1%